OpenStudy (anonymous):

plz help If the first term of a geometric sequence is 300 and the third term is 3, find the sum of the first 25 terms.

OpenStudy (anonymous):

u can use the formula which is An=ar^(n-1) then when n=1 we have A1=ar^0=a=100 when n=3, we have A3=ar^2=3,since a=100, then 100r^2=3, r=(3/100)^0.5 then for sum of first 25, when is sum from n=1 to n=25, u can use the formula here: http://en.wikipedia.org/wiki/Geometric_progression just plug n=25, r=(3/100)^0.5, a=100 into the formula then u ll get the answer, hope it helps